A researcher records the following scores on a working memory quiz for two samples. Which sample has the laegest standard deviation? Sample A: 2,3,4,5,6,7 and 8
Sample B: 4,5,6,7,8,9, and 10
Answer #1

Sample A : 2,3,4,5,6,7,8 mean = EX-35 - 5 Sample Stendard deviation (5) = [(x=* $ = (2-5)*+--- +(8-5) nol 7-1 = 14.66667 = 2.=54666667 = 52.16027 deviation of booth the sample Hence, i8 the Standard Same. W